OpenStudy (anonymous):

The point-slope form of the equation of the line that passes through (–9, –2) and (1, 3) is y – 3 = 1/2 (x – 1) . What is the slope-intercept form of the equation for this line? a. y=1/2x + 2 b. y=1/2x - 4 c. y= 1/2x + 5/2 d. y=1/2x - 7/2

OpenStudy (anonymous):

Just first multiply 2 with left hand side term and and show me your work.

OpenStudy (anonymous):

The given equation is \[y-3=1/2(x-1)\] Open the bracket on the right side \[y-3=(x/2)-(1/2)\] take the 3 to the right side \[y=(x/2) +3 -(1/2)\] simplifies to \[y=(x/2)+(5/2)\] which is of the form \[y=mx+c\] \[m=1/2\] \[c=5/2\] option C